Solve the simultaneous equation +=5 and +=1


1
Expert's answer
2021-06-22T08:49:00-0400

"\displaystyle
a_1x + b_1y = 5\\
a_2x + b_2y = 1\\

a_1a_2x + b_1a_2y = 5a_2\\
a_1a_2x + a_1b_2y = a_1\\

(b_1a_2 - a_1b_2)y = 5a_2 - a_1\\

y = \frac{5a_2 - a_1}{b_1a_2 - a_1b_2}\\

a_1b_2x + b_1b_2y = 5b_2\\
b_1a_2x + b_1b_2y = b_1\\

(a_1b_2 - b_1a_2)x = 5b_2 - b_1\\

x = \frac{5b_2 - b_1}{a_1b_2 - b_1a_2}\\"